# The direct linearization scheme with the Lamé function: The KP equation and reductions

### v2

@@ 3.2.2 Marchenko equation
Similar to the treatment in section 2.2, here we again restrict our discussion to the case where $x,g_{2},g_{3}\in\mathbb{R}$, $w_{1}>0$, $\mathrm{Im}w_{2}>0$ and $\Delta>0$. These settings allow us to have an alternative elliptic DL scheme which can yield real non-singular solutions for the KdV equation.
@@
$u(x,t)=-2\wp(x+w_{2})-2\partial_{x}\int_{L}\varphi(x,t;l)\varrho_{l}(t)\widetilde{\Psi}_{x}(l)\mathrm{d}\lambda(l).$ (3.11)
@@
In [18], the Marchenko equation (3.14) was directly introduced by assuming $F(x,s,t)$ satisfies a kind of Lax pair of the KdV equation ( see equations (5) and (6) in [18]). As discussed in section 2.2, when $k$ and $l$ take values within the interval $(0,w_{1})$, the validity of integral (3.13) can be ensured. At this point, the integration contour should be adjusted to encircle the discrete points $k_{m}\in(0,w_{1})$. Consequently, $K(x,s,t)$ and $F(x,s,t)$ are written as
@@ 3.2.3 Real-valued elliptic solitons
We assume that $L$ is a contour in the fundamental rectangular domain with vertices $-w_{1}-w_{2},w_{1}-w_{2},w_{1}+w_{2},-w_{1}+w_{2}$, $\{k_{j}\}$ belongs to the domain circled by $L$, and the measure $\mathrm{d}\lambda(l)$ is taken as
@@
$u=-2\wp(x)-2\partial_{x}(\mathbf{r}^{T}\mathbf{u}^{(0)})=-2\wp(x)+2(\ln\tau)_{xx},$ (3.20)
@@
$\tau=\tau_{N}=\sum_{J\subset S}\left(\prod_{i\in J}c_{i}\right)\Biggl{(}\prod_{i,j\in J\atop i<j}A_{i,j}\Biggr{)}\frac{\sigma(x+2\sum_{i\in J}k_{i}+w_{2})}{\sigma(x+w_{2})\prod_{i\in J}\sigma(2k_{i})}\exp\left(\sum_{i\in J}\theta_{[e]}(x,t;k_{i})\right),$ (3.24)
@@
$c_{i}=\widetilde{c}_{i}\exp\left(-2\zeta(w_{2})k_{i}\right),~{}~{}~{}\widetilde{c}_{i}\in\mathbb{R}.$ (3.25)
@@
$\tau_{1}=1+c_{1}\frac{\sigma(x+w_{2}+2k_{1})}{\sigma(x+w_{2})\sigma(2k_{1})}e^{\theta_{[e]}(x,t;k_{1})},$ (3.26)
@@
$\tau_{2}$ $=1+c_{1}\frac{\sigma(x+w_{2}+2k_{1})}{\sigma(x+w_{2})\sigma(2k_{1})}e^{\theta_{[e]}(x,t;k_{1})}+c_{2}\frac{\sigma(x+w_{2}+2k_{2})}{\sigma(x+w_{2})\sigma(2k_{2})}e^{\theta_{[e]}(x,t;k_{2})}$
$+c_{1}c_{2}A_{1,2}\frac{\sigma(x+w_{2}+2k_{1}+2k_{2})}{\sigma(x+w_{2})\sigma(2k_{1})\sigma(2k_{2})}e^{\theta_{[e]}(x,t;k_{1})+\theta_{[e]}(x,t;k_{2})}.$ (3.27)
@@ 3.3.2 Marchenko equation
For the Boussinesq equation, to obtain real-valued elliptic soliton solutions, we should require all three elliptic cube roots to be real, which means that $\wp^{\prime}(z)=\wp^{\prime}(k)$ should have three distinct real roots within the real period for a given $k\in\mathbb{R}$. However, according to Proposition B.1, in the case where $\Delta>0$, there are no distinct real roots for $\wp^{\prime}(z)=\wp^{\prime}(k)$ in $(0,2w_{1}]$. We exclude the case $\Delta<0$ where distinct real roots exist, because prefer the formulation of $w_{1}$ being real and $w_{2}$ purely imaginary so that we can shift the singularity to the complex plane. Thus, we need to modify some settings to reduce the Marchenko equation to the Boussinesq case.
@@
allows three distinct real roots $z=\omega_{j}(l)$, $j=0,1,2$ in $(0,2w_{1}]$ when $g_{2},g_{3}\in\mathbb{R}$ and $\Delta>0$ (see Appendix B.2). In the following we always set $\omega_{j}(l)=l$. Choosing the measure
@@
$u(X,Y)=-2\wp(X+w_{2})-2\partial_{X}\sum_{j=1}^{2}\int_{L_{j}}\psi(X,Y;l)\gamma_{-\omega_{j}(l)}(Y)\widetilde{\Psi}_{X}(-\omega_{j}(l))\mathrm{d}\lambda_{j}(l).$ (3.39)
@@ 3.3.3 Real-valued elliptic solitons
$u(X,Y)=-2\wp(X+w_{2})-2(\mathbf{s}^{T}(\mathbf{I+M})^{-1}\mathbf{r})_{X}=-2\wp(X+w_{2})+2(\ln\tau)_{XX},$ (3.50)
@@
$\theta_{[e]}(X,Y;k_{i},\omega_{\alpha}(k_{i}))=(-\zeta(k_{i})+\zeta(\omega_{\alpha}(k_{i})))X+(\wp(k_{i})-\wp(\omega_{\alpha}(k_{i})))Y-\zeta(w_{2})(k_{i}-\omega_{\alpha}(k_{i})),$
$F_{J}=\frac{\sigma(X+w_{2}+\sum_{i\in J}(k_{i}-\sum_{\alpha=1}^{2}\nu_{i\alpha}\omega_{\alpha}(k_{i})))}{\sigma(X+w_{2})\prod_{i\in J}\sigma(k_{i}-\sum_{\alpha=1}^{2}\nu_{i\alpha}\omega_{\alpha}(k_{i}))},$
@@
$u_{1}(X,Y)=-2\wp(X+w_{2})+2\bigg{(}\ln\Big{(}1+\sum_{\alpha=1}^{2}c_{1\alpha}\frac{\sigma(X+k_{1}-\omega_{\alpha}(k_{1})+w_{2})}{\sigma(X+w_{2})\sigma(k_{1}-\omega_{\alpha}(k_{1}))}e^{\theta_{[e]}(X,Y;k_{1},\omega_{\alpha}(k_{1}))}\Big{)}\bigg{)}_{XX}.$ (3.54)
@@
$\tau=1+c_{11}\frac{\sigma(X+k_{1}-\omega_{1}(k_{1})+w_{2})}{\sigma(X+w_{2})\sigma(k_{1}-\omega_{1}(k_{1}))}e^{\theta_{[e]}(X,Y;k_{1},\omega_{1}(k_{1}))}$ (3.56)
@@
$\tau=1+c_{11}\frac{\sigma(X+k_{1}-\omega_{1}(k_{1})+w_{2})}{\sigma(X+w_{2})\sigma(k_{11}-\omega_{1}(k_{1}))}e^{\theta_{[e]}(X,Y;k_{1},\omega_{1}(k_{1}))}$
$\quad+c_{21}\frac{\sigma(X+k_{2}-\omega_{1}(k_{2})+w_{2})}{\sigma(X+w_{2})\sigma(k_{2}-\omega_{1}(k_{2}))}e^{\theta_{[e]}(X,Y;k_{2},\omega_{1}(k_{2}))}+c_{11}c_{21}\frac{\sigma(k_{1}-k_{2})\sigma(\omega_{1}(k_{1})-\omega_{1}(k_{2}))}{\sigma(k_{1}-\omega_{1}(k_{2}))\sigma(\omega_{1}(k_{1})-k_{2})}$
×$\quad\times\frac{\sigma(X+k_{1}-\omega_{1}(k_{1})+k_{2}-\omega_{1}(k_{2})+w_{2})}{\sigma(X+w_{2})\sigma(k_{1}-\omega_{1}(k_{1}))\sigma(k_{2}-\omega_{1}(k_{2}))}e^{\theta_{[e]}(X,Y;k_{1},\omega_{1}(k_{1}))+\theta_{[e]}(X,Y;k_{2},\omega_{1}(k_{2}))}.$ (3.57)
@@ 4 Conclusions
In this paper, we have established the elliptic DL schemes and Marchenko equations for the KP, KdV and Boussinesq equations. Non-singular real elliptic soliton solutions for these equations were obtained. The DL scheme of the KP equation consists of an integral equation (e.g.(2.9)) and a formula of solution (e.g.(2.12)). This is confirmed by checking the Lax pair. Since the Marchenko equation involves integration on the real axis, we have discussed real-valuedness of the Weierstrass functions in Appendix A. Based on these discussions, we can choose parameters $k$ and $l^{\prime}$ such that the integration (2.35) on $(x,\infty)$ makes sense. Then we were able to construct a Marchenko equation for elliptic soliton solutions (see (2.37)). To have non-singular real solutions, in addition to utilizing real-valuedness of the Weierstrass functions, we also employed a trick, $x\to x+w_{2}$, to shift zeros of $\sigma(z)$ from the real axis to the half complex plane. As a result, we obtained $\tau$ function for non-singular real elliptic solitons of the KP equation. The illustration shows line solitons on a periodic background. We also obtained resonant Y-shape elliptic line solitons.
@@ Appendix A Real-valuedness of the Weierstrass functions
where $e_{i}=\wp(w_{i})$, and $w_{3}=w_{1}+w_{2}$. Non-zero discriminant $\Delta=g_{2}^{3}-27g_{3}^{2}$ ensures the smoothness of the curve. The other two Weierstrass functions, $\zeta(z)$ and $\sigma(z)$, are quasi-periodic, see (2.7).
@@
where ‘bar’ denotes complex conjugate, and $f$ can be either Weierstrass-$\wp$, $\zeta$ or $\sigma$. When $z\in\mathbb{R}$, we immediately prove the proposition.
@@
(1) Fundamental half periods $w_{1}$ is real, $w_{2}$ is purely imaginary. $w_{1}=a+\mathrm{i}b$ and $w_{2}=a-\mathrm{i}b$ ($a>0$, $b>0$).
(2) Values at half-periods $e_{1}>0\geq e_{3}>e_{2}$; $\zeta(w_{1})>0$, $\mathrm{i}\zeta(w_{2})\in\mathbb{R}$; $e_{1}^{2}>\frac{g_{2}}{12}$, $e_{2}^{2}>\frac{g_{2}}{12}$.
@@
$e_{3}\geq 0$, $e_{1}=-\alpha+\mathrm{i}\beta$, $e_{2}=\overline{e_{1}}$, ($\alpha\geq 0$, $\beta>0$); $\zeta(w_{3})>0$, $\mathrm{i}\zeta(w_{1}-w_{2})\in\mathbb{R}$; $e_{3}^{2}>\frac{g_{2}}{12}$;
@@
$z\in(0,2w_{3})$, $\wp(z)\geq e_{3}$;
$z\in(0,2w_{1}-2w_{2})$, $\wp(z)\leq e_{3}$.
(4) Real-value of $\Psi_{z}(a)$ $z\in\mathbb{R}$ and $a\in(0,2w_{1})$, $(w_{2},w_{2}+2w_{1})$, (modulo the period lattice). $z\in\mathbb{R}$ and $a\in(0,2w_{3})$, (modulo the period lattice).
@@
$-\wp(z;g_{2},g_{3})=\wp(\mathrm{i}z;g_{2},-g_{3}),\qquad-\mathrm{i}\zeta(z;g_{2},g_{3})=\zeta(\mathrm{i}z;g_{2},-g_{3}).$ (A.3)
When $g_{3}<0$, we denote the half-periods of $\wp(z;g_{2},g_{3})$ by $\widetilde{w}_{j}$, and $\tilde{e}_{j}=\wp(\tilde{w}_{j};g_{2},g_{3})$, $j=1,2,3$, where $\tilde{w}_{3}=\tilde{w}_{1}+\tilde{w}_{2}$. When $\Delta>0$, one can take $\tilde{w}_{1}=-\mathrm{i}w_{2}>0$, $\tilde{w}_{2}=\mathrm{i}w_{1}$ (see page 105 of [4]), thus we have
$\tilde{e}_{1}>\tilde{e}_{3}>0>\tilde{e}_{2},\quad\tilde{e}^{2}_{1}>\frac{g_{2}}{12},\quad\tilde{e}^{2}_{2}>\frac{g_{2}}{12},\quad\zeta(\tilde{w}_{1};g_{2},g_{3})\in\mathbb{R},\quad\mathrm{i}\zeta(\tilde{w}_{2};g_{2},g_{3})>0.$ (A.4)
When $\Delta<0$, one may take $\tilde{w}_{1}=\mathrm{i}w_{2}$, $\tilde{w}_{2}=-\mathrm{i}w_{1}$, thus we have
$\tilde{e}_{3}<0,\quad\tilde{e}^{2}_{3}>\frac{g_{2}}{12},\quad\zeta(\tilde{w}_{3};g_{2},g_{3})\in\mathbb{R},\quad\mathrm{i}\zeta(\tilde{w}_{1}-\tilde{w}_{2};g_{2},g_{3})>0.$ (A.5)

### v3

@@ 3.2.2 Marchenko equation
Similar to the treatment in section 2.2, here we again restrict our discussion to the case where $x,g_{2},g_{3}\in\mathbb{R}$, $w>0$, $\mathrm{Im}w^{\prime}>0$ and $\Delta>0$. These settings allow us to have an alternative elliptic DL scheme which can yield real non-singular solutions for the KdV equation.
@@
$u(x,t)=-2\wp(x+w^{\prime})-2\partial_{x}\int_{L}\varphi(x,t;l)\varrho_{l}(t)\widetilde{\Psi}_{x}(l)\mathrm{d}\lambda(l).$ (3.11)
@@
In [30], the Marchenko equation (3.14) was directly introduced by assuming $F(x,s,t)$ satisfies a kind of Lax pair of the KdV equation ( see equations (5) and (6) in [30]). As discussed in section 2.2, when $k$ and $l$ take values within the interval $(0,w)$, the validity of integral (3.13) can be ensured. At this point, the integration contour should be adjusted to encircle the discrete points $k_{m}\in(0,w)$. Consequently, $K(x,s,t)$ and $F(x,s,t)$ are written as
@@ 3.2.3 Real-valued elliptic solitons
We assume that $L$ is a contour in the fundamental rectangular domain with vertices $w-w^{\prime},w-w^{\prime},w+w^{\prime},-w+w^{\prime}$, $\{k_{j}\}$ belongs to the domain circled by $L$, and the measure $\mathrm{d}\lambda(l)$ is taken as
@@
$u=-2\wp(x)-2\partial_{x}(\mathbf{r}^{T}\mathbf{u}^{(0)})=2\partial_{x}^{2}\ln\big{(}\sigma(x)\tau\big{)},$ (3.20)
@@
$\tau=\tau_{N}=\sum_{J\subset S}\left(\prod_{i\in J}c_{i}\right)\Biggl{(}\prod_{i,j\in J\atop i<j}A_{i,j}\Biggr{)}\frac{\sigma(x+2\sum_{i\in J}k_{i}+w^{\prime})}{\sigma(x+w^{\prime})\prod_{i\in J}\sigma(2k_{i})}\exp\left(\sum_{i\in J}\theta_{[e]}(x,t;k_{i})\right),$ (3.24)
@@
$c_{i}=\widetilde{c}_{i}\exp\left(-2\zeta(w^{\prime})k_{i}\right),~~~\widetilde{c}_{i}\in\mathbb{R}.$ (3.25)
@@
$\tau_{1}=1+c_{1}\frac{\sigma(x+w^{\prime}+2k_{1})}{\sigma(x+w^{\prime})\sigma(2k_{1})}e^{\theta_{[e]}(x,t;k_{1})},$ (3.26)
@@
$\tau_{2}$ $=1+c_{1}\frac{\sigma(x+w^{\prime}+2k_{1})}{\sigma(x+w^{\prime})\sigma(2k_{1})}e^{\theta_{[e]}(x,t;k_{1})}+c_{2}\frac{\sigma(x+w^{\prime}+2k_{2})}{\sigma(x+w^{\prime})\sigma(2k_{2})}e^{\theta_{[e]}(x,t;k_{2})}$
$+c_{1}c_{2}A_{1,2}\frac{\sigma(x+w^{\prime}+2k_{1}+2k_{2})}{\sigma(x+w^{\prime})\sigma(2k_{1})\sigma(2k_{2})}e^{\theta_{[e]}(x,t;k_{1})+\theta_{[e]}(x,t;k_{2})}.$ (3.27)
@@ 3.3.2 Marchenko equation
For the Boussinesq equation, to obtain real-valued elliptic soliton solutions, we should require all three elliptic cube roots to be real, which means that $\wp^{\prime}(z)=\wp^{\prime}(k)$ should have three distinct real roots within the real period for a given $k\in\mathbb{R}$. However, according to Proposition B.1, in the case where $\Delta>0$, there are no distinct real roots for $\wp^{\prime}(z)=\wp^{\prime}(k)$ in $(0,2w]$. We exclude the case $\Delta<0$ where distinct real roots exist, because prefer the formulation of $w$ being real and $w^{\prime}$ purely imaginary so that we can shift the singularity to the complex plane. Thus, we need to modify some settings to reduce the Marchenko equation to the Boussinesq case.
@@
allows three distinct real roots $z=\omega_{j}(l)$, $j=0,1,2$ in $(0,2w]$ when $g_{2},g_{3}\in\mathbb{R}$ and $\Delta>0$ (see Appendix B.2). In the following we always set $\omega_{j}(l)=l$. Choosing the measure
@@
$u(X,Y)=-2\wp(X+w^{\prime})-2\partial_{X}\sum_{j=1}^{2}\int_{L_{j}}\psi(X,Y;l)\gamma_{-\omega_{j}(l)}(Y)\widetilde{\Psi}_{X}(-\omega_{j}(l))\mathrm{d}\lambda_{j}(l).$ (3.39)
@@ 3.3.3 Real-valued elliptic solitons
$u(X,Y)=-2\wp(X+w^{\prime})-2(\mathbf{s}^{T}(\mathbf{I+M})^{-1}\mathbf{r})_{X}=2\partial_{X}^{2}\ln\big{(}\sigma(X+w^{\prime})\tau\big{)},$ (3.50)
@@
$\theta_{[e]}(X,Y;k_{i},\omega_{\alpha}(k_{i}))=(-\zeta(k_{i})+\zeta(\omega_{\alpha}(k_{i})))X+(\wp(k_{i})-\wp(\omega_{\alpha}(k_{i})))Y-\zeta(w^{\prime})(k_{i}-\omega_{\alpha}(k_{i})),$
$F_{J}=\frac{\sigma(X+w^{\prime}+\sum_{i\in J}(k_{i}-\sum_{\alpha=1}^{2}\nu_{i\alpha}\omega_{\alpha}(k_{i})))}{\sigma(X+w^{\prime})\prod_{i\in J}\sigma(k_{i}-\sum_{\alpha=1}^{2}\nu_{i\alpha}\omega_{\alpha}(k_{i}))},$
@@
$u_{1}(X,Y)=2\partial_{X}^{2}\ln\Big{(}\sigma(X+w^{\prime})+\sum_{\alpha=1}^{2}c_{1\alpha}\frac{\sigma(X+k_{1}-\omega_{\alpha}(k_{1})+w^{\prime})}{\sigma(k_{1}-\omega_{\alpha}(k_{1}))}e^{\theta_{[e]}(X,Y;k_{1},\omega_{\alpha}(k_{1}))}\Big{)}.$ (3.54)
@@
$\tau=1+c_{11}\frac{\sigma(X+k_{1}-\omega_{1}(k_{1})+w^{\prime})}{\sigma(X+w^{\prime})\sigma(k_{1}-\omega_{1}(k_{1}))}e^{\theta_{[e]}(X,Y;k_{1},\omega_{1}(k_{1}))}$ (3.56)
@@
$\tau=1+c_{11}\frac{\sigma(X+k_{1}-\omega_{1}(k_{1})+w^{\prime})}{\sigma(X+w^{\prime})\sigma(k_{11}-\omega_{1}(k_{1}))}e^{\theta_{[e]}(X,Y;k_{1},\omega_{1}(k_{1}))}$
$\quad+c_{21}\frac{\sigma(X+k_{2}-\omega_{1}(k_{2})+w^{\prime})}{\sigma(X+w^{\prime})\sigma(k_{2}-\omega_{1}(k_{2}))}e^{\theta_{[e]}(X,Y;k_{2},\omega_{1}(k_{2}))}+c_{11}c_{21}\frac{\sigma(k_{1}-k_{2})\sigma(\omega_{1}(k_{1})-\omega_{1}(k_{2}))}{\sigma(k_{1}-\omega_{1}(k_{2}))\sigma(\omega_{1}(k_{1})-k_{2})}$
$\quad\times\frac{\sigma(X+k_{1}-\omega_{1}(k_{1})+k_{2}-\omega_{1}(k_{2})+w^{\prime})}{\sigma(X+w^{\prime})\sigma(k_{1}-\omega_{1}(k_{1}))\sigma(k_{2}-\omega_{1}(k_{2}))}e^{\theta_{[e]}(X,Y;k_{1},\omega_{1}(k_{1}))+\theta_{[e]}(X,Y;k_{2},\omega_{1}(k_{2}))}.$ (3.57)
@@ 4 Conclusions
In this paper, we have established the elliptic DL schemes and Marchenko equations for the KP, KdV and Boussinesq equations. Non-singular real elliptic soliton solutions for these equations were obtained. The DL scheme of the KP equation consists of an integral equation (e.g.(2.9)) and a formula of solution (e.g.(2.12)). This is confirmed by checking the Lax pair. Since the Marchenko equation involves integration on the real axis, we have discussed real-valuedness of the Weierstrass functions in Appendix A. Based on these discussions, we can choose parameters $k$ and $l^{\prime}$ such that the integration (2.35) on $(x,\infty)$ makes sense. Then we were able to construct a Marchenko equation for elliptic soliton solutions (see (2.37)). To have non-singular real solutions, in addition to utilizing real-valuedness of the Weierstrass functions, we also employed a trick, $x\to x+w^{\prime}$, to shift zeros of $\sigma(z)$ from the real axis to the half complex plane. As a result, we obtained $\tau$ function for non-singular real elliptic solitons of the KP equation. The illustration shows line solitons on a periodic background. We also obtained resonant Y-shape elliptic line solitons.
@@ Appendix A Real-valuedness of the Weierstrass functions
where $e_{i}=\wp(w_{i}),i=1,2,3,$ and $w_{1}=w$, $w_{2}=w+w^{\prime}$, $w_{3}=w^{\prime}$. Non-zero discriminant $\Delta=g_{2}^{3}-27g_{3}^{2}$ ensures the smoothness of the curve. The other two Weierstrass functions, $\zeta(z)$ and $\sigma(z)$, are quasi-periodic, see (2.7).
@@
where ‘bar’ denotes complex conjugate, and $f$ can be either $\wp$, $\zeta$ or $\sigma$. When $z\in\mathbb{R}$, we immediately prove the proposition.
@@
(1) Fundamental half periods $w_{1}$ is real, $w_{3}$ is purely imaginary. $w_{1}=a+\mathrm{i}b$ and $w_{3}=a-\mathrm{i}b$ ($a>0$, $b>0$).
(2) Values at half-periods $e_{1}>0\geq e_{2}>e_{3}$; $\zeta(w_{1})>0$, $\mathrm{i}\zeta(w_{3})\in\mathbb{R}$; $e_{1}^{2}>\frac{g_{2}}{12}$, $e_{3}^{2}>\frac{g_{2}}{12}$.
@@
$e_{2}\geq 0$, $e_{1}=-\alpha+\mathrm{i}\beta$, $e_{3}=\overline{e_{1}}$, ($\alpha\geq 0$, $\beta>0$); $\zeta(w_{2})>0$, $\mathrm{i}\zeta(w_{3}-w_{1})\in\mathbb{R}$; $e_{2}^{2}>\frac{g_{2}}{12}$;
@@
$z\in(0,2w_{2})$, $\wp(z)\geq e_{2}$;
$z\in(0,2w_{1}-2w_{3})$, $\wp(z)\leq e_{2}$.
(4) Real-value of $\Psi_{z}(a)$ $z\in\mathbb{R}$ and $a\in(0,2w_{1})$, $(w_{3},w_{3}+2w_{1})$, (modulo the period lattice). $z\in\mathbb{R}$ and $a\in(0,2w_{2})$, (modulo the period lattice).
@@
$\wp(z;g_{2},g_{3})=\wp(\mathrm{i}z;g_{2},-g_{3}),\qquad-\mathrm{i}\zeta(z;g_{2},g_{3})=\zeta(\mathrm{i}z;g_{2},-g_{3}).$ (A.3)
When $g_{3}<0$, we denote the half-periods of $\wp(z;g_{2},g_{3})$ by $\widetilde{w}_{j}\,(j=1,3)$, and $\tilde{e}_{j}=\wp(\tilde{w}_{j};g_{2},g_{3})\,(j=1,2,3)$, where $\tilde{w}_{2}=\tilde{w}_{1}+\tilde{w}_{3}$. When $\Delta>0$, one can take $\tilde{w}_{1}=-\mathrm{i}w_{3}>0$, $\tilde{w}_{3}=\mathrm{i}w_{1}$ (see page 105 of [5]), thus we have
$\tilde{e}_{1}>\tilde{e}_{2}>0>\tilde{e}_{3},\quad\tilde{e}^{2}_{1}>\frac{g_{2}}{12},\quad\tilde{e}^{2}_{3}>\frac{g_{2}}{12},\quad\zeta(\tilde{w}_{1};g_{2},g_{3})\in\mathbb{R},\quad\mathrm{i}\zeta(\tilde{w}_{3};g_{2},g_{3})>0.$ (A.4)
When $\Delta<0$, one may take $\tilde{w}_{1}=\mathrm{i}w_{3}$, $\tilde{w}_{3}=-\mathrm{i}w_{1}$, thus we have
$\tilde{e}_{2}<0,\quad\tilde{e}^{2}_{2}>\frac{g_{2}}{12},\quad\zeta(\tilde{w}_{2};g_{2},g_{3})\in\mathbb{R},\quad\mathrm{i}\zeta(\tilde{w}_{1}-\tilde{w}_{3};g_{2},g_{3})>0.$ (A.5)